201212 R1ProgressRep

From New IAC Wiki
Jump to navigation Jump to search


Sector 1 - Summary of Results:

- ADB continues to have problems with noise, and today we were unable to get it to reliably show a cosmic pulse. Switched to just using the post-amp for now. Removed the grounding strap from the S1 Chamber. No noticeable difference in noise levels.

- During cosmic pulse counting, Mac determined the new nominal voltage for sense wires should be 1400V to avoid a saturation condition that results in a large characteristic pulse consistently showing up.

- Mac also suggested we start keeping a log of current levels whenever running HV testing, and ensure the over-current trip settings are correct on the HV power supply (0.1 sec, 1μA)

- J-lab should be sending us some new equipment to serve the role of the ADB crate - either a new ADB or some DCRB's.

Sector 5 - strung 51 Sense wires (now on SL1 - row 200).

Sector 6 - tested SL2 through row 154. The decision was made with Mac, George, Dr. Forest and Dr. McNulty that post-potting testing will not continue after today. The post-potting Tension Trend data will be compared with pre-potting Tension Trend data to check for any unexpected results.


We only had two stringers today, so we focused on finishing the Sense wire stringing in Sector 5.

Sector 5 - strung 81 Sense wires. All stringing now complete. Will begin visual inspection and final pre-potting testing tomorrow.


Sector 1 - Dr. Forest and David continued testing. The following is from David's report on today's progress:

"Pulse Generator seems to be functioning right now.

Proceeded to gather counting efficiency data for various HV settings on the wires. Discovered that during ramp-up the sense wires exceeded the 1µA safety cut-off. Raised the cut-off to 3 µA and the sense wires were able to ramp-up to 1400V without the safety tripping. Current stabilized at 0.115 µA on sense, 0.190 µA on field, and 0.000 µA on guards after ramp-up.

A significant amount of noise is generated in the post-amp signal after turning on the DC power supply for the STB, peak to peak amplitude averages around 50 mV. Turned up the discriminator's threshold to 100 mV for an initial test run. Experimenting with the trigger level on the oscilloscope, I found that 75 mV is a "safe" level to use for triggering only on cosmic pulses and not on noise. Even at the 75 mV threshold, there are occasional noise pulses that exceed that level and cause false counts (pulse does not conform to what a normal cosmic looks like). Consulted with Dr. Forest, he thinks a couple of the noise pulses we're seeing are from the HVAC system (confirmed this by turning off the HV, and noise pulses still came through the post-amp), and some others may be spark discharges on the ground-line from other equipment in the building. He suggested running longer count times (10+ min.) to make the noise pulses a negligible part of the count.

I took counts running from 600-1,000 sec at a few different voltage settings, and we got results that are in line with what you'd expect - upping the voltage results in a higher count rate. I already took one background sampling for 50 minutes over lunch break, and we are taking another overnight to get a larger sample of background while no one is likely to be working in the building in other labs." (David Coneff 12/5/12).

Sector 5 - tested Sense wires strung on Tuesday (row 200-226). Began visual inspection. Began SL2 Field pre-potting testing. Tested row 2-40 (267 Field wires).

Completed mylar window installation supply inventory for Sectors 2,3,5 and 6. Will compile notes and e-mail George a final list of supplies we have and need.


Sector 5 - pre-potting testing SL2 Fields row 41-184 (1008 Field wires). Completed weekly Tension Trend test.

Completed weekly cleanroom clean.


Sector 1 - soldered jumpers on 7 SL1 HV boards.

Sector 5 - pre-potting testing SL2 Fields row 185-204 and SL1 Fields row 185-191 (189 total wires).


Sector 1 - continued testing high voltage plateau and confirmed multiple hits using the ADB crate.

Sector 5 - pre-potting testing - completed SL2 Fields. Tested SL1 Fields row 3-48. (483 total wires).

Received our particle counter from Kanomax after calibration. Particle readings from this afternoon on will be taken with our counter.


Sector 1 - soldered jumpers on 5 HV boards.

David continued testing. DAQ is now setup with the post-amp and discriminators to function on 16 channels. Threshold for discriminators is now set to a nominal -1.0 V to eliminate noise (HVAC system, grounding noise, etc.). Double counts are still occasionally happening on a given channel even with the higher threshold. Determined the DAQ can handle up to a 12 KHz event rate. Started testing all 16 channels to confirm signal injection works reliably with the DAQ

Sector 5 - pre-potting testing SL1 Fields row 49-114 (462 total wires).


Sector 1 - Soldered jumpers on 3 HV boards. Began setting up scintillators to begin coincidence counting of cosmic rays.

Sector 5 - pre-potting testing Sl1 Fields row 115-197 (531 total wires). Completed weekly Tension Trend test.

Completed weekly cleanroom clean.


Sector 5 - pre-potting testing SL2 Sense row 3-80 and SL1 Sense row 3-70 (432 total wires).


Sector 5 - pre-potting testing SL2 Sense row 81-226 and SL1 Sense row 71-226 (906 total wires).


High temperature alarm. Decreased temperature set point on HVAC by one degree. Temperatures are now within stringing specifications.

Sector 5 - visual inspection and replacement of wires that are out of specifications.